# The Inertial Spin Model of flocking with position-dependent forces

### v2

@@ III Numerical results
To gain some insight into the behavior of the modified model, we consider the case of a single particle in an external field to study the confinement effects. We solve the equations of motion numerically in their second-order form (20). We have employed an integration scheme used in Brownian Dynamics simulations Rapaport (2004), based on a Velocity Verlet integrator with Lagrange multipliers to enforce fixed speed. The procedure is described in detail in ref. Cavagna et al. (2016). We consider the simplest form of the force to achieve confinement, namely a simple harmonic force
@@
The mean squared displacement (MSD) $d^{2}(t)=[\mathbf{x}(t)-\mathbf{x}(0)]^{2}$ is shown in Fig. 1 for the ISM with two different values of the dimensionless parameter $\Omega$ (23a) (recall $\Omega\gg 1$ means inertia dominates, while $\Omega\ll 1$ corresponds to overdamped systems) and different field strengths (i.e. different $k_{0}$). In all cases, for very short times there is a ballistic regime ($d^{2}\sim t^{2}$) which crosses over to a diffusive ($d^{2}\sim t$) regime for the free case. If the confining field is present, the displacement eventually saturates at an $\Omega$-dependent plateau (Fig. 2). For weak enough fields, the diffusive regime can be observed before the plateau, but at high fields the confinement effects manifest before the transition from ballistic to diffusive motion.
@@
Figure 2: Maximum (plateau) of the mean squared displacement vs. $\Omega$ for the ISM with harmonic confinement, $k_{0}=1$.
@@
The effects of confinement can also be noticeable in the velocity correlation function $C_{v}(t)=\langle\mathbf{v}(t)\cdot\mathbf{v}(0)\rangle$ (Fig. 3). If the confining field strength is large enough, there is a range of times for which anticorrelation is observed and followed, at high $\Omega$, by damped oscillations. The first anti-correlation minimum should not be mistaken for an inertial effect: it is due to the fact that the velocity changes direction due to the confinement force, and it is actually observed both at low and high $\Omega$ for large enough $k_{0}$. For $k_{0}=1$ it can be seen that $C_{v}(t)$ has its first minimum around $t\approx 3$, which is the time at which confinement effects start to be noticeable in $d^{2}(t)$ (cf. top panels of Fig. 1). At this time, the diffusive regime has not yet been reached, and correlation function of the free particle is still non-vanishing. In contrast, for weaker fields (e.g. $k_{0}=0.02$ in Fig. 3) it can happen that the MSD starts deviating towards the plateau already in the diffusive regime, when the velocity has lost correlation (as measured by $C_{v}(t)$ for the free case). In this case no anticorrelation minimum is observed. Inertial effects manifest, for large $k_{0}$, in the presence of damped oscillations in $C_{v}(t)$, lasting roughly during the transition from the ballistic regime to the plateau in the MSD. For all $k_{0}$, when $\Omega$ is large the correlation function has a flat derivative as $t\to 0$, a sign of second-order dynamics (see discussion in Cavagna et al. (2017)). This can be seen plotting $h(x)=-\log C(x)/x$ with $x=t/\tau$ and $\tau$ is the correlation time (Fig. 4). $h(x)$ tends to a constant for $x\to 0$ if $C(t)$ resembles a simple exponential for short times (overdamped system), or to 0 if $C(t)$ has a flat derivative. We have computed $\tau$ using the spectral definition Halperin and Hohenberg (1969),
@@
The velocity time correlations look similar to those of an harmonic oscillator with Langevin dynamics, a system that might be considered a minimal model for a single confined particle. However, the behavior of the active system is different from the harmonic oscillator’s. One way to see this is considering the correlation times of the three dynamical variables position $\mathbf{x}$, velocity $\mathbf{v}$ and spin $\mathbf{s}$. Since the spin can be computed from the trajectories as $\mathbf{s}=\mu\mathbf{v}\times\dot{\mathbf{v}}$, this definition can also be applied to trajectories of the harmonic oscillator to obtain a correlation function $C_{s}(t)$, so we can compute the time correlation functions for position $C_{x}(t)$, velocity $C_{v}(t)$ and spin $C_{s}(t)$ in both systems. Writing the oscillator equation as
$m\ddot{\mathbf{x}}+\eta\dot{\mathbf{x}}+k\mathbf{x}=\xi(t)$ (29)
@@
the inertial parameter is $\Omega_{\text{HO}}=km/\eta^{2}$. The ratios $\tau_{v}/\tau_{x}$ and $\tau_{s}/\tau_{x}$ show quite different behavior as a function of $\Omega$ or $\Omega_{\text{HO}}$ (Fig. 5). In the HO, both ratios increase monotonically as the system becomes more underdamped, while in the ISM $\tau_{v}/\tau_{x}$ is monotonically decreasing and $\tau_{s}/\tau_{x}$ reaches a maximum near $\Omega=1$ and then decreases. The third ratio, $\tau_{s}/\tau_{v}$, is monotonically increasing in both cases, but as $\Omega\to 0$ it tends to 1 for the HO while it goes to zero for the ISM, as in the latter the spin correlation time vanishes more quickly than that of the velocity.
@@
Figure 5: Correlation times ratios vs. inertial parameter for the 2-$d$ stochastic harmonic oscillator (left) and ISM (right).
@@
It is harder to discriminate between ISM and HO working at fixed $\Omega$ (which would be the situation in experimental observations). But in principle one could use $h(x)$ for the velocity correlation function $C_{v}(t)$ to estimate whether $\Omega$ is high or low, then study the relaxation time ratios. If $h(x)$ stays finite for $x\to 0$ (low $\Omega$), then $\tau_{s}/\tau_{v}<1$ is not compatible with simple harmonic motion, and neither is the presence of a ballistic regime in $d^{2}(t)$. At high inertia instead, $\tau_{s}/\tau_{x}$ and $\tau_{v}/\tau_{x}$ both less than one and of similar value would be incompatible with an HO.
@@
Finally, we have considered a measure of the trajectories’ shape. The sample trajectories in Fig. 1 show that in the inertial case the particle finds it more difficult to turn back on itself, and responds to confinement making turns with smoother curvature than in the over-damped case. This suggests that measuring the number of times the trajectory intersects with itself might be a way to detect inertial effects directly from the trajectory. We show the number of times the trajectory intersects itself during one velocity correlation time in Fig. 6. It is clear that at low inertia the trajectory crosses itself much more often than at high $\Omega$. This tendency is also observed in the harmonic oscillator, the ISM case there seems to be a rather sharp crossover around $\Omega=1$.
Figure 6: Number of times a trajectory crosses itself during one velocity correlation time vs. inertial parameter. Top: ISM, bottom: 2-$d$ stochastic harmonic oscillator. Insets show, as an example, the trajectories for $\Omega=1$ for each system, with circles indicating the intersections.
@@ IV Discussion and conclusions
We have shown how the inertial spin model equations of motion must be modified to include position-dependent forces in this model. The most compact formulation is in terms of velocity and spin, Eqs. (II.2), where the positional forces enter only in the equation for the spin.
As a simple application of the new equations, we have considered a single harmonically confined active particle, and shown how inertia can be detected in this simple case without collective effects. The appearance of an anti-correlation minimum in the velocity temporal self-correlation function is an effect of strong confinement and it appears in over- and under-damped systems, but longer-lasting damped oscillations are only found with high inertia. Such damped oscillations were recently reported for male Anopheles gambiae (malaria mosquitoes) in laboratory swarms Cavagna et al. (2023b). Also, the trajectories recorded in that work look more “open” than the Vicsek trajectories, displaying loops roughly similar to the simulated trajectories at high $\Omega$ in Fig. 1. The present developments are useful for attempting to apply the ISM to the analysis of these and other experiments on laboratory-confined swarms (e.g. Ni et al. (2015); Gorbonos et al. (2016)).
@@
As a next step, it is clearly of interest to study systems of many particles with open boundary conditions, maintaining a finite density through confinement or cohesive interactions. This can be done with the present formulation of the inertial spin model, and it is worth pursuing to achieve a better description of experimental results as well as to gain further theoretical understanding of the collective properties of active models. As an example, in a series of recent papers González-Albaladejo et al. (2023); González-Albaladejo and Bonilla (2023a, b), it has been claimed that, for the Vicsek model, replacing periodic boundaries with an harmonic confinement alters the behavior of the model near ordering, giving rise to a phase transition characterized by scale-free chaos and an extended criticality region and yielding different static and dynamic critical exponents. Open conditions thus deserve deeper inquiry, both for over- and under-damped inertial systems. Finally, position-dependent forces can also be used to investigate the response of a swarm to external perturbations that do not directly alter the velocities. Work in these directions is in progress.

### v3

@@ III Numerical results
To gain some insight into the behavior of the modified model, we consider the case of a single particle in an external field to study the confinement effects. We solve the equations of motion numerically in their second-order form (20). We have employed an integration scheme used for the ISM in ref. Cavagna et al. (2016), which has been used in Brownian Dynamics simulations Rapaport (2004) and based on a Velocity Verlet integrator with Lagrange multipliers to enforce fixed speed (see Appendix A for details). We consider the simplest form of the force to achieve confinement, namely a simple harmonic force
@@
This is the simplest situation one can think of beyond a free particle. We choose this as the first application of the modified ISM because we can clearly gauge the effects of the position-dependent field, which is the new element in the model, and check that it effectively produces the desired effect, i.e. confinement. But this scenario is also useful for the case of swarms of insects in laboratory conditions, in closed environments (e.g. Puckett et al. (2014); Ni et al. (2015); Gorbonos et al. (2016); Cavagna et al. (2023b)) and low densities Reynolds (2023), where interactions can be mostly neglected. We ask whether inertial effects can be inferred from the trajectory of a single individual, and whether in this situation the ISM can be distinguished from other simple models of a confined particle. Thus we will compare results at high and very low $\Omega$ (essentially equivalent to comparing the ISM and Vicsek models), and also to two other simple models of confined inertial particles: the Brownian (massive) harmonic oscillator and the active Brownian particle Bechinger et al. (2016); Zöttl and Stark (2023).
The Brownian harmonic oscillator (HO) is a useful comparison, partly in view of Eq. (25), but mostly as a null model, being the simplest non-active model for a confined, fluctuating particle. Clearly the complications of active models are not worth if the resulting phenomenology cannot be distinguished from that of the HO. The Browinan HO is described by the inertial Langevin equation
$m\ddot{\mathbf{x}}+\eta\dot{\mathbf{x}}+k_{0}\mathbf{x}=\xi(t),$ (28)
and the inertial parameter describing the underdamped–overdamped crossover is $\Omega_{\text{HO}}=k_{0}m/\eta^{2}$. The dynamic equations for the inertial active Brownian particle (ABP) Enculescu and Stark (2011) with an harmonic confining field are, in 2-$d$
$m\ddot{\mathbf{x}}$ $=-\eta\dot{\mathbf{x}}-k_{0}\mathbf{x}+\eta v_{0}(\cos\theta,\sin\theta)+\xi(t),$ (29)
$\dot{\theta}$ $=\zeta,$ (30)
where $\theta(t)$ is the heading direction, $v_{0}$ will be held constant and $\xi(t)$ and $\zeta(t)$ are Gaussian white noise processes with zero mean and variances $2T\eta$ and $2D_{r}$ respectively ($T$ is the temperature and $D_{r}$ is a rotational diffusion constant). The inertial parameter is $\Omega_{\text{ABP}}=k_{0}m/\eta^{2}$.
@@
A quantity that must be sensitive to confinement effects is the mean squared displacement (MSD) $d^{2}(t)=[\mathbf{x}(t)-\mathbf{x}(0)]^{2}$, which we show in Fig. 2 for the ISM with two different values of the dimensionless parameter $\Omega$ (23a) (recall $\Omega\gg 1$ means inertia dominates, while $\Omega\ll 1$ corresponds to overdamped systems) and different field strengths (i.e. different $k_{0}$). The MSD curves are obtained as an average over 100 trajectories, after discarding an initial time so that the MSD is stationary. In all cases, for very short times there is a ballistic regime ($d^{2}\sim t^{2}$) which crosses over to a diffusive ($d^{2}\sim t$) regime for the free case. If the confining field is present, the displacement eventually saturates at an $\Omega$-dependent plateau (Fig. 3). For weak enough fields, the diffusive regime can be observed before the plateau, but at high fields the confinement effects manifest before the transition from ballistic to diffusive motion. The presence of the external field brings a qualitative change in the shape of $d^{2}(t)$ in the expected direction, i.e. a plateau appears indicating the particle is confined. However, there is no qualitative difference between high and low $\Omega$. The ballistic regime results from the persitence of the velocity (i.e. the fact that velocities have a finite correlation time, see below). In the usual Brownian motion, persistence is a result of an inertial term, but the pure Vicsek particle (the overdamped $\Omega\to 0$ limit) also leads to a persistent random walk at low enough temperature, so all the curves of Fig. 2 look qualitatively the same (except for the obvious fact that there is no plateau if $k_{0}=0$). Increasing $\Omega$ delays the onset of the diffusive regime, which is hardly observable when both $\Omega$ and $k_{0}$ are high. Also, the value of the plateau itself depends on $\Omega$ and on $k_{0}$ (Fig. 3). The asymptotic $d^{2}$ goes as $\sim 1/k_{0}$, similarly to the Brownian harmonic oscillator where by equipartition one has $\langle x^{2}\rangle=k_{B}T/k_{0}$. On the other hand, the plateau grows (almost linearly in the explored range) with $\Omega$, quite differently from the Browinan oscillator, where the confinement scale is independent of the mass. The ABP (not shown) behaves very similar to the ISM (the plateau goes as $\sim 1/k_{0}$ and depends on $\Omega_{\text{ABP}}$), except that the plateau dependence with inertia is clearly superlinear for $\Omega_{\text{ABP}}$ greater than approximately 2.
@@
Figure 3: Dependence of the plateau of the mean squared displacement on $\Omega$ and confinement strength $k_{0}$ for the harmonically confined ISM.
@@
The effects of confinement can also be noticeable in the velocity correlation function $C_{v}(t)=\langle\mathbf{v}(t)\cdot\mathbf{v}(0)\rangle$ (Fig. 4). If the confining field strength is large enough, there is a range of times for which anticorrelation is observed and followed, at high $\Omega$, by damped oscillations. The first anti-correlation minimum should not be mistaken for an inertial effect: it is due to the fact that the velocity changes direction due to the confinement force, and it is actually observed both at low and high $\Omega$ for large enough $k_{0}$. For $k_{0}=1$ it can be seen that $C_{v}(t)$ has its first minimum around $t\approx 3$, which is the time at which confinement effects start to be noticeable in $d^{2}(t)$ (cf. top panels of Fig. 2). At this time, the diffusive regime has not yet been reached, and the correlation function of the free particle is still non-vanishing. In contrast, for weaker fields (e.g. $k_{0}=0.02$ in Fig. 4) it can happen that the MSD starts deviating towards the plateau already in the diffusive regime, when the velocity has lost correlation (as measured by $C_{v}(t)$ for the free case). In this case no anticorrelation minimum is observed. Inertial effects manifest, for large $k_{0}$, in the presence of damped oscillations in $C_{v}(t)$, lasting roughly during the transition from the ballistic regime to the plateau in the MSD. For all $k_{0}$, when $\Omega$ is large the correlation function has a flat derivative as $t\to 0$, a sign of second-order dynamics (see discussion in Cavagna et al. (2017)). This can be seen plotting $h(x)=-\log C(x)/x$ with $x=t/\tau$ and $\tau$ is the correlation time (Fig. 5). $h(x)$ tends to a constant for $x\to 0$ if $C(t)$ resembles a simple exponential for short times (overdamped system), or to 0 if $C(t)$ has a flat derivative. For the correlation time $\tau$ we have used the spectral definition Halperin and Hohenberg (1969),
@@
The extrapolation of $h(x)$ to $x=0$ is appreciably different from 0 only for rather low values of $\Omega$, dropping rather sharply to very low values around $\Omega=10^{-1}$ (see Fig. 7).
@@
The velocity time correlations look similar to those of an harmonic oscillator with Langevin dynamics. However, the behavior of the active system is different from the harmonic oscillator’s. One way to see this is considering the correlation times of the three dynamical variables position $\mathbf{x}$, velocity $\mathbf{v}$ and spin $\mathbf{s}$. Since the spin can be computed from the trajectories as $\mathbf{s}=\mu\mathbf{v}\times\dot{\mathbf{v}}$, this definition can also be applied to trajectories of the harmonic oscillator and the ABP, even if the spin does not play a significant role in those models. So we can compute the time correlation functions for position $C_{x}(t)$, velocity $C_{v}(t)$ and spin $C_{s}(t)$ in the three systems. The ratios $\tau_{v}/\tau_{x}$ and $\tau_{s}/\tau_{x}$ show quite different behavior as a function of $\Omega$, $\Omega_{\text{HO}}$ or $\Omega_{\text{ABP}}$ (Fig. 6). In the HO, both ratios increase monotonically as the system becomes more underdamped, while in the ISM $\tau_{v}/\tau_{x}$ is monotonically decreasing and $\tau_{s}/\tau_{x}$ reaches a maximum near $\Omega=1$ and then decreases. The third ratio, $\tau_{s}/\tau_{v}$, is monotonically increasing in both cases, but as $\Omega\to 0$ it tends to 1 for the HO while it goes to zero for the ISM, as in the latter the spin correlation time vanishes more quickly than that of the velocity. In the ABP, $\tau_{s}$ and $\tau_{v}$ are always very similar and are both smaller than $\tau_{x}$. Their value relative to $\tau_{x}$ increases monotonically with $\Omega_{\text{ABP}}$, and as $\Omega_{\text{ABP}}$ grows all correlation times tend to coincide.
@@
Figure 6: Correlation times ratios vs. inertial parameter for the ISM (top), 2-$d$ stochastic harmonic oscillator (middle) and active Browinian particle (bottom).
@@
If the relaxation times could be computed at several values of the inertia, it would be possible to discriminate among the three models. However, this is harder when working at fixed $\Omega$ (which would be the typical situation in experimental observations). But in principle one could use $h(x)$ for the velocity correlation function $C_{v}(t)$ to estimate whether $\Omega$ is high or low, then study the relaxation time ratios. If $h(x)$ stays finite for $x\to 0$ (low $\Omega$), then $\tau_{s}/\tau_{v}<1$ is not compatible with simple harmonic motion, and neither is the presence of a ballistic regime in $d^{2}(t)$. At high inertia instead, $\tau_{s}/\tau_{x}$ and $\tau_{v}/\tau_{x}$ both less than one and of similar value would be incompatible with an HO. Also $\tau_{s}/\tau_{v}$ much different from 1 would exclude the ABP, at any value of inertia.
@@
Finally, we have considered a measure of the trajectories’ shape. The sample trajectories in Fig. 2 show that in the inertial case the particle finds it more difficult to turn back on itself, and responds to confinement making turns with smoother curvature than in the over-damped case. This suggests that measuring the number of times the trajectory intersects with itself might be a way to detect inertial effects directly from the trajectory. We show the number of self-intersections during one velocity correlation time in Fig. 7. It is clear that at low inertia the trajectory crosses itself much more often than at high $\Omega$. This tendency is also observed in the harmonic oscillator and the ABP. A rather sharp drop is observed at a model-dependent value of $\Omega$. Inspecting the trajectories directly, it is clear that this number decreases as trajectories turn less on themselves and tend to make smoother loops rather than sharp turns. However, it is not evident that there is a clear qualitative change in their shapes across the crossover. Fig. 7 also shows the value $h_{0}$ of $h(x)$ extrapolated to $x=0$. Although it also displays a rather sharp crossover, it occurs at a value of $\Omega$ about a hundred times smaller than that of the crossings. The two quantities thus reflect different properties of the trajectories.
Figure 7: Number of times a trajectory crosses itself during one velocity correlation time vs. inertial parameter for the ISM, harmonic oscillator and ABP (middle panel). The values have been normalized to 1 at $\Omega=0$ to be able to plot them together, since the value at small $\Omega$ varies much for the three models. Bottom panel: $h(x)$ (see Fig. 5) extrapolated to $x=0$ for the ISM. Top panel: example trajectories near the crossover in crossings.
@@ IV Discussion and conclusions
We have shown how the inertial spin model equations of motion must be modified to include position-dependent forces. The most compact formulation is in terms of velocity and spin, Eqs. (18), where the positional forces enter only in the equation for the spin. The ISM was proposed as an extension of the Vicsek model, and thus it introduces activity in the same way, which is perhaps the simplest one can imagine: by imposing a hard constraint on the particles’ speed. This, plus the ISM’s own Hamiltonian structure, greatly limits the way in which position-dependent forces may be added to this model. Another approach to introducing activity is to endow the particles with a “heading” or “spin”, which is an internal degree of freedom related to but different from the velocity, as in active Brownian particles Bechinger et al. (2016); Zöttl and Stark (2023). This is also the approach of the purely Hamiltonian, conservative, active model of Bore et al. (2016). Interestingly, in that model the Hamiltonian structure also places limits in the kind of interaction term one can consider (in that case the interaction between velocity and spin Casiulis et al. (2020)).
As a simple application of the new equations, we have considered a single harmonically confined active particle, and shown how inertia can be detected in this simple case without collective effects. We have shown the confined ISM shows behavior qualitatively different from both the Brownian harmonic oscillator (our null, non-active, model for confined motion) and the active Brownian particle, another simple model for self-propelled particles. The three models show a plateau in the mean-square displacement that depends in the same way on the strength of the confining field (measured by the value of the harmonic constant $k_{0}$). However, the active models distinguish themselves from the harmonic oscillator in that the plateau also depends on the value of the inertial parameter $\Omega$ or $\Omega_{\text{ABP}}$. Up to the explored values of $\Omega$, this dependence is linear for the ISM, while a non-linear regime is evident in the ABP.
We have also studied the time correlation functions that can be computed from the trajectory. The effect of strong confinement is the appearance of an anti-correlation minimum in the velocity temporal self-correlation. This anti-correlation appears when the velocity is forced to change direction at times shorter than the correlation time of the free particle under otherwise identical parameters; it appears in over- and under-damped systems and is unrelated to inertia. When present, inertia manifests in the slope of the correlation function at short times (Fig. 5), and in the presence of damped oscillations beyond the first anti-correlation minimum. Such damped oscillations were recently reported for male Anopheles gambiae (malaria mosquitoes) in laboratory swarms Cavagna et al. (2023b). Also, the trajectories recorded in that work appear to show relatively smooth loops, similar to the simulated trajectories at high $\Omega$ (Fig. 2), rather than the more tortuous trajectories of the low-$\Omega$, Vicsek-like case. Thus, inertia seems an important ingredient to include in the modelling, even for the study of single individuals, making the present developments useful for attempting to apply the ISM to the analysis of these and other experiments on laboratory-confined swarms (e.g. Ni et al. (2015); Gorbonos et al. (2016)). However, the velocity correlation function alone is not enough to distinguish among the harmonic oscillator, the ISM and the ABP, as the different parameters in each model can be tuned to give very similar correlation functions. We have shown that the models display clear differences when one examines the trends of the relaxation times and the plateau in the mean squared displacement (Figs. 3 and 6). Unfortunately in the experimental case one will usually not be able to tune $\Omega$, so discriminating among models will be harder, but some conclusions can be drawn examining together the MSD and the three relaxation times.
@@
As a next step, it is clearly of interest to study systems of many particles with open boundary conditions, maintaining a finite density through confinement or cohesive interactions. This can be done with the present formulation of the inertial spin model, and it is worth pursuing to achieve a better description of experimental results as well as to gain further theoretical understanding of the collective properties of active models. As an example, in a series of recent papers González-Albaladejo et al. (2023); González-Albaladejo and Bonilla (2023a, b), it has been claimed that, for the Vicsek model, replacing periodic boundaries with an harmonic confinement alters the behavior of the model near ordering, giving rise to a phase transition characterized by scale-free chaos and an extended criticality region and yielding different static and dynamic critical exponents. Open conditions thus deserve deeper inquiry, both for over- and under-damped inertial systems. Finally, position-dependent forces can also be used to investigate the response of a swarm to external perturbations that do not directly alter the velocities.
In summary, we have shown how positional forces should be added to the Inertial Spin Model, preserving its Hamiltonian structure, and presented evidence that a single confined particle displays qualitatively different behavior in the ISM than in two other simple models of confinement (one equilibrium, one active). This development makes it possible to use the ISM without periodic boundary conditions, by introducing either a confining field or cohesion. The addition of these new ingredients allows to apply the ISM to a broader range of conditions, and to simulate more closely various aspects of biological groups.
@@ Acknowledgements.
Appendix A Details of the numerical simulations
To simulate the ISM or the Brownian harmonic oscillator we integrate numerically the corresponding differential equations. In the case of the ISM model, since we have the constraint $\mathbf{v}^{2}(t)=v_{0}$ we use an integration scheme used in Brownian Dynamics which allows for an exact implementation of said constraint via Lagrange multipliers, and which in the underdamped case $\Omega\to\infty$ reduces to the velocity Verlet integrator widely used in Molecular Dynamics due to its good energy conservation properties and computational affordability Rapaport (2004). The only drawbacks is that the overdamped case with $\chi$ strictly equal to 0 cannot be integrated using this method. This scheme has been employed before in simulations of the ISM, e.g. in Ref. Cavagna et al. (2016).
To arrive at an integration algorithm we start from the second order equation (20), which we rewrite as
$\frac{d^{2}\mathbf{v}}{dt^{2}}=\frac{v_{0}^{2}}{\chi}\left[\mathbf{F}+\mathbf{F}_{v}+\mathbf{f}_{c}\right]$ (32)
where the first term represents the position-dependent external (in our particular simulation case $\mathbf{F}=-\frac{k_{0}}{v_{0}}\mathbf{x}$), the next term includes the random and viscous forces and $\mathbf{f}_{c}$ is the constraint force. At the continuous level it is given by the rest of (20), but for the numerical integration it is computed as explained below, so that the constraint is enforced exactly. To obtain the discretized equations we integrate (32) assuming $F$ varies linearly with time in a small interval $\Delta t$. The term $\mathbf{f}_{c}$ is disregarded at first and later reintroduced as explained below. Defining $\mathbf{a}=\frac{d\mathbf{v}}{dt}$, $\mathbf{b}=\frac{d\mathbf{a}}{dt}$ one arrives at
$\mathbf{r}(t+\Delta t)=\mathbf{r}(t)+\Delta t\mathbf{v}(t)$ (33)
$\begin{split}&\mathbf{v}(t+\Delta t)=\mathbf{v}(t)+\Delta tc_{1}\mathbf{a}(t)+(\Delta t)^{2}c_{2}\mathbf{b}(t)\\ &+(\Delta t)^{2}c_{2}\zeta_{1}\mathbf{v}(t)+\Theta_{v}(t)\end{split}$ (34)
$\begin{split}&\mathbf{a}(t+\Delta t)=c_{0}\mathbf{a}(t)+(c_{1}-c_{2})\Delta t\left[\mathbf{b}(t)+\zeta_{1}\mathbf{v}(t)\right]+\\ &c_{2}\Delta t\left[\mathbf{b}(t+\Delta t)+\zeta_{2}\mathbf{v}(t+\Delta t)\right]+\Theta_{a}(t)\end{split}$ (35)
$\mathbf{b}(t+\Delta t)=\frac{v_{0}^{2}}{\chi}\mathbf{F}(\mathbf{r}(t+\Delta t))$ (36)
where $\lambda$ and $\mu$ are related to the constraint, the other constants result from the integration as
$c_{0}=e^{-\eta\Delta t/\chi}$ (37)
$c_{1}=\frac{\chi}{\eta\Delta t}(1-c_{0})$ (38)
$c_{2}=\frac{\chi}{\eta\Delta t}(1-c_{1}),$ (39)
and $\Theta_{v}$ and $\Theta_{a}$ are random variables related to the random force. They are independent for each axis and each pair of components is drawn form a bivariate Gaussian distribution with zero first moments and second moments given by
$\Theta_{v}^{2}\right>$ $=\frac{T\chi}{\eta^{2}}\left(2\frac{\eta}{\chi}\Delta t-3+4c_{0}-c_{0}^{2}\right)$ (40)
$\Theta_{a}^{2}\right>$ $=\frac{T}{\chi}\left(1-c_{0}^{2}\right)$ (41)
$\Theta_{v}\Theta_{a}\right>$ $=\frac{T}{\eta}\left(1-c_{0}\right).$ (42)
The discrete equations reduce to the velocity Verlet integrator for Molecular Dynamics Allen and Tildesley (1987); Swope et al. (1982) in the underdamped $\eta\rightarrow 0$ limit.
The constraint is enforced as in the RATTLE algorithm Andersen (1983), only that since the constraints on each particle are independent, the Lagrange multipliers can be found analytically and there is no need of an iterative procedure. Imposing $\mathbf{v}^{2}(t+\Delta t)=v_{0}^{2}$ and $\mathbf{v}(t+\Delta t)\cdot\mathbf{a}(t+\Delta t)=0$ one obtains
$\zeta_{1}$ $=\frac{\omega_{+}-1}{(\Delta t)^{2}c_{2}},$ (43)
$\zeta_{2}$ $=-\frac{\mathbf{v}(t+\Delta t)\cdot\mathbf{a}^{\prime}(t+\Delta t)}{c_{2}v_{0}^{2}\Delta t},$ (44)
where $\omega_{+}$ is the positive root of
$v_{0}^{2}\omega^{2}+2\mathbf{v}\cdot\Delta\mathbf{v}\omega+\Delta v^{2}=v_{0}^{2},$ (45)
$\Delta\mathbf{v}=c_{1}\Delta t\mathbf{a}(t)+c_{2}(\Delta t)^{2}\mathbf{b}(t),$ (46)
and $\mathbf{a}^{\prime}(t+\Delta t)$ is equal to $\mathbf{a}(t+\Delta t)$ without the term proportional to $\zeta_{2}$.
Each step is performed in two stages, as in the velocity Verlet scheme Allen and Tildesley (1987). First the random variables are drawn, $\mathbf{r}$ is updated, $\mathbf{a}$ is partially updated using only the terms that depend on quantities evaluated at $t$, $\mathbf{v}$ is updated and then the constraint terms are computed and applied. Second the forces for the new positions and velocities are computed, and finally the update of $\mathbf{a}$ is completed.
For the results reported here, we have set $\Delta t=0.001$ and the friction coefficient and temperature respectively to $\eta=1$ and $T=1$.